Flash Calculations

A feed stream enters a flash unit at 60 ^{o}C. The composition of the stream as well as the Antoine coefficients are given below:

Compound  Flow rate (kmol/hr) B.Pt.(^{o}C) A  B  C
Hexane  20  68.7  6.87776 1171.530 224.366
1,2 Dichloroethane 30  83.4  7.04636 1305.400 230.000
Methylcyclohexane 50  100  6.82689 1272.864 221.630

Antoine Equation:

log_{10}(P) = A − \frac{B}{T + C}

where P is the vapor pressure in bars and T is the temperature in ^{o} C.

1. The flash operates at 80 ^{o}C and 1 bar. What is the composition of the liquid and vapor streams exiting the flash unit?

2. The flash operates at 80 ^{o}C and it is desired to have a vapor fraction of 20 % (\frac{V}{F} = 0.20). What is the pressure at which the flash unit operates and what is the composition of the liquid and vapor streams exiting the flash unit?

3. The flash operates at 1 bar and it is desired to have a vapor fraction of 25 % (\frac{V}{F} = 0.25). What is the temperature at which the flash unit operates and what is the composition of the liquid and vapor streams exiting the flash unit?

4. The flash operates at 80 ^{o}C and it is desired that the split fraction of 1,2 Dichloroethane in the vapor phase be 0.25 (ξ_{2} = 0.25). Compute the pressure at which the flash unit operates and the composition of the liquid and vapor streams exiting the flash unit.

Instructions

• Do the above problems in CHEMCAD.

• How do the results change if you use different “K-value” methods? Tabulate your results to show this comparison.

• Do these problems by “hand calculations”.